OpenStudy (anonymous):

PERCENT OF CHANGE? 295 pounds to 245 Pounds

OpenStudy (anonymous):

actual change is 295-245 = 50 percent change is \[50\div 295= =.1694=17\% \] rounded

OpenStudy (anonymous):

295 - 245 = 50 50/295 = 0.1694915254237288 0.1694915254237288 * 100 = 16.94915254237288% Proof: 0.1694915254237288 * 295 = 50.

OpenStudy (anonymous):

50/295 = 0.1695 = 16.95%

OpenStudy (anonymous):

rounded

OpenStudy (anonymous):

Percent of change p = Amount of Increase or decrease divide by the Original Amount. 295-245 = 50 17%

OpenStudy (anonymous):

17% decrease, or -17%

OpenStudy (anonymous):

Rounded to what place? Nearest ones? Nearest tenth? Nearest hundreth?